# Stronger ViTs With Octic Equivariance

### v2

@@ Abstract
Recent efforts at scaling computer vision models have established Vision Transformers (ViTs) as the leading architecture. ViTs incorporate weight sharing over image patches as an important inductive bias. In this work, we show that ViTs benefit from incorporating equivariance under the octic group, i.e., reflections and 90-degree rotations, as a further inductive bias. We develop new architectures, octic ViTs, that use octic-equivariant layers and put them to the test on both supervised and self-supervised learning. Through extensive experiments on DeiT-III and DINOv2 training on ImageNet-1K, we show that octic ViTs yield more computationally efficient networks while also improving performance. In particular, we achieve approximately 40% reduction in FLOPs for ViT-H while simultaneously improving both classification and segmentation results.
@@ 1 Introduction
In the pursuit of flexible yet scalable models, Vision Transformers (ViTs) [21] have emerged as the dominant architecture in modern computer vision. Key to their success is the combination of visual tokens, constructed from image patches, with the powerful attention mechanism [48], resulting in an extremely versatile yet scalable architecture. This scalability is due in large part to weight-sharing between the tokens, which ensures permutation equivariance.
Equivariance provides a powerful inductive bias in neural networks by enforcing structured responses to transformations such as translations, rotations or reflections. Another major benefit of equivariance is the reduction in computational cost from weight-sharing due to equivariance constraints. It is therefore interesting to investigate whether imbuing models with further equivariance yields stronger performance and better scaling. To this end, equivariance to roto-reflections, formalized through the octic group $\mathrm{D}_{8}$, is a particularly attractive group, as many vision tasks exhibit such symmetries, making it a natural inductive bias.
$\mathrm{D}_{8}$ equivariance was introduced to convolutional networks by Cohen and Welling [13], demonstrating improved parameter efficiency by constraining the learnable space through group structure [55, 3, 52]. Yet, despite its theoretical appeal, state-of-the-art vision models [39, 37, 31, 51] do not incorporate such structure. We argue that this is not due to a lack of utility, but to practical limitations: most existing implementations construct equivariant layers using computationally inefficient architectures, leading to increased FLOPs and runtime. Without efficient, hardware-compatible implementations, these methods have remained impractical for large-scale models. As a result, the potential of equivariant design remains largely unexplored in state-of-the-art systems, in particular in the context of ViTs. Recently, Bökman et al. [8] showed that equipping ViTs with horizontal flip equivariance results in retained performance while saving FLOPs. However, their improvements are limited due to the small cardinality of their chosen group.
In this paper, we demonstrate that scaling equivariance to larger groups can be efficiently implemented – yielding faster, stronger, and more compact models, cf. Figure 1. Specifically, we introduce octic-equivariant layers for ViTs, leveraging the $\mathrm{D}_{8}$ symmetry group of $90^{\circ}$ rotations and reflections, cf. Figure 2. Our approach integrates seamlessly into existing ViT architectures and leads to significant gains in throughput and memory efficiency, without sacrificing accuracy.
@@
    We introduce octic-equivariant layers for ViTs, described in Section 3, that, when incorporated, achieve substantially higher FLOP efficiency given their representational capacity.
@@
    We develop an efficient implementation, including a custom Triton kernel for the activation function, to enable fast and scalable octic-equivariant layers.
@@
    We validate our approach through extensive experiments in Section 4 on two representative ViT models (DeiT III and DINOv2), showcasing strong performance gains at scale.
Figure 1: Computational savings. Using octic layers in ViTs significantly reduces the computational complexity while maintaining or improving classification accuracy on ImageNet-1K, for both supervised and self-supervised training. Detailed results can be found in Section 4.
@@ 2 Related Work
Vision Transformer.
The Vision Transformer (ViT) was introduced by Dosovitskiy et al. [21] and has subsequently achieved state-of-the-art results in many domains of computer vision [10, 39, 31, 22, 50]. Zhai et al. [57] showed that ViTs benefit from scaling up to 2B parameters and Dehghani et al. [17] extended this to 22B parameters. To support scaling, Chen et al. [12] explored sparse ViTs, while techniques from the language domain—such as parallel layers [49]—have been adopted to improve efficiency. Additionally, Alabdulmohsin et al. [1] investigated which architectural dimensions are most effective to scale. In this work, we propose the incorporation of octic layers in ViTs to increase efficiency at scale. Our motivation follows the likes of Swin Transformers [34] and SparseViT [12], making use of inherent properties of images to improve efficiency.
@@
The equivariance of CNNs to (cyclic) image translations can be extended to incorporate larger symmetry groups such as rotations and reflections, as shown by Cohen and Welling [13], Dieleman et al. [20] using Group Equivariant CNNs (G-CNNs). Cohen and Welling [14], Weiler and Cesa [52] generalized G-CNNs to steerable CNNs, where the features transform according to general group representations. Our octic ViTs can be seen as a more scalable ViT analogue of the octic steerable CNNs by Cohen and Welling [14]. There have also been prior efforts on attention- and Transformer-based equivariant vision architectures [41, 56, 33]. Our work is part of an ongoing research direction of studying and improving the scalability of equivariant networks [6, 4, 5, 36, 47]. Prior work in this direction mostly focused on point cloud data, with the notable exception of Bökman et al. [8] who considered images. In particular, Bökman et al. [8] demonstrated that incorporating horizontal mirroring equivariance into modern image classifiers increases compute efficiency while maintaining representational power. Building on this, we extend the symmetry to the octic group and show in more extensive experiments the practical usefulness of equivariance for two popular ViT training objectives.
Figure 2: Architecture. Patches are first extracted from an image using specialized octic filters and the resulting features are processed by $k$ octic ViT blocks. The final patch embeddings can be fed to $l-k$ standard Transformer blocks (as demonstrated by our $\mathcal{H}_{8}$ and $\mathcal{I}_{8}$ ViTs) or be used for, e.g., classification by producing an invariant quantity, as in our $\mathrm{D}_{8}$ (ViT).
@@ 3 Method
Group equivariant networks are networks that respect a data symmetry [55, 13]. For surveys of equivariant networks we refer to [7, 25, 53], and for background on group and representation theory we refer to [44]. In this section, we design octic-equivariant ViT layers.
We begin by preliminaries for octic equivariance in Section 3.1, followed by specifics of the Transformer layers in Section 3.2 and a detailed discussion of computational efficiency in Section 3.3. We summarize the most important notation in Section 3.1.1.
@@ 3.1 Preliminaries on Octic Equivariance
In this work we focus on the dihedral group with eight elements, $\mathrm{D}_{8}$ =$=\{e,r,r^{2},r^{3},s,sr,sr^{2},sr^{3}\}$, such that $r^{4}=s^{2}=e$ is the identity element and $r^{3}=srs$ [^1] The reader is cautioned that $\mathrm{D}_{8}$ is sometimes alternatively denoted $\mathrm{D}_{4}$. We think of $\mathrm{D}_{8}$ as acting on images by reflections $s$ and 90°90°90  ° rotations $r$, as illustrated in Figure 3. $\mathrm{D}_{8}$ is also called the octic group, and we opt for this shorter name throughout.
@@
$\begin{split}&\rho_{\mathrm{A1}}(r)=\rho_{\mathrm{A1}}(s)=1;\qquad\quad\rho_{\mathrm{A2}}(r)=1,\ \rho_{\mathrm{A2}}(s)=-1;\\ &\rho_{\mathrm{B1}}(r)=-1,\ \rho_{\mathrm{B1}}(s)=1;\quad\rho_{\mathrm{B2}}(r)=\rho_{\mathrm{B2}}(s)=-1;\\ &\text{and}\quad\rho_{\mathrm{E}}(r)=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\ \rho_{\mathrm{E}}(s)=\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}.\end{split}$ (1)
We use the same notation as the original work on steerable CNNs [14] for these irreps, but choose a different basis for $\rho_{E}$. It is known from elementary representation theory that any representation of $\mathrm{D}_{8}$ can be decomposed into irreps as
@@
so that $\rho_{\text{reg}}(g)$ can be written as a permutation matrix. Importantly, $\rho_{\text{reg}}$ commutes with pointwise activation functions such as GELU [27].
@@
The regular representation $\rho_{\text{reg}}$ can be block-diagonalized to its isotypical decomposition $\rho_{\text{iso}}$ through (2) as $\rho_{\text{reg}}(g)=Q_{\text{reg}}\rho_{\text{iso}}(g)Q_{\text{reg}}^{-1}$ with
@@
It is a general fact for finite groups that the regular representation decomposes into all irreps with multiplicities equal to their dimensions. The change of basis $Q_{\text{reg}}$ (written out in full in Appendix A) is the inverse Fourier transform of $\mathrm{D}_{8}$, with $Q_{\text{reg}}^{-1}=Q_{\text{reg}}^{\mathsf{T}}$ being the Fourier transform.
(a) PatchEmbed filters.
Figure 3: (a) PatchEmbed filters from a trained network. More filters are shown in the appendix, Figure 6.
(b) Octic action on patchified images.
(c) Octic action on $\rho_{\text{iso}}$-features.
Equivariant networks use equivariant linear layers, which map between representations. A linear $G$-equivariant map or intertwiner, $W\in\mathbb{R}^{n\times n}$, between representations $\rho_{1},\rho_{2}$, commutes with their action, i.e., $\rho_{2}(g)W=W\rho_{1}(g)$. It follows from Schur’s lemma that $W=\lambda I$ for some scalar $\lambda$ if $\rho_{1},\rho_{2}$ are irreps, and that $\lambda=0$ if $\rho_{1},\rho_{2}$ are not isomorphic [44, Section I.2.2]222 We can apply Schur’s lemma for complex representations here since the irreps listed in Example 3.1 are irreducible over the complex numbers. We will however only use real-valued linear layers, i.e. $\lambda\in\mathbb{R}$.. As $\rho_{\text{iso}}$ is just a stack of irreducible representations, any intertwiner between such representations will be sparse, which is why they are less computationally expensive than ordinary linear layers. The naive computational complexity for a linear map $W:\mathbb{R}^{|\mathrm{D}_{8}|}\to\mathbb{R}^{|\mathrm{D}_{8}|}$ is $|\mathrm{D}_{8}|^{2}=64$ operations. In contrast, intertwiners $\rho_{\text{iso}}\to\rho_{\text{iso}}$ require a total of $\sum_{i}^{k}m_{i}^{2}d_{i}=1+1+1+1+2^{2}\cdot 2=12$ operations for $\mathrm{D}_{8}$. From a signal processing perspective, this is analogous to the computational saving of convolution being point-wise multiplications in frequency space.
@@
Square images can be considered as elements of $\mathbb{R}^{M\times M\times C}$ where $M$ is the image height/width in pixels and $C$ is the number of colour channels (typically $C=3$). There is a natural $\mathrm{D}_{8}$-representation $\rho_{\text{image}}$ associated with square images, where $\rho_{\text{image}}(r)$ is the pixel permutation rotating the images anti-clockwise by 90°90°90  ° and $\rho_{\text{image}}(s)$ is the permutation reflecting the images left-to-right.
@@
In ViTs, features are elements of $\mathbb{R}^{N\times N\times D}$, which we will think of as $N^{2}\times D$ matrices. Here, $N$ is the number of image tokens along the height/width of the image, so $N=M/P$ where $P$ is the patch size (typically $P=14$ or $P=16$) and $D$ is the channel dimension. The simplest representation that we consider on features is the permutation representation $\rho_{\text{token}}$ that, analogously to $\rho_{\text{image}}$, permutes the tokens according to elements of $\mathrm{D}_{8}$.
@@
We can equip the channel dimension of ViT features with a group representation $\rho_{\text{chan}}$ to obtain “steerable” features. If $D$ is divisible by $8$, we can consider multiples of $\rho_{\text{reg}}$ or $\rho_{\text{iso}}$ as $\rho_{\text{chan}}$. The complete representation $\rho$ acting on the $N^{2}\times D$-matrix $\mathbf{x}$, is then permuting the tokens according to $\rho_{\text{token}}$ and modifying the channels according to $\rho_{\text{chan}}$. Concretely,
$\rho(g)\mathrm{Vec}(\mathbf{x})=\mathrm{Vec}\left(\rho_{\text{token}}(g)\mathbf{x}\rho_{\text{chan}}(g)^{\mathsf{T}}\right)=\left(\rho_{\text{chan}}(g)\otimes\rho_{\text{token}}(g)\right)\mathrm{Vec}(\mathbf{x}),$ (6)
where $\mathrm{Vec}(\mathbf{x})$ is the column-wise vectorization of the matrix $\mathbf{x}$ and $\otimes$ is the tensor product of representations or (equivalently) the Kronecker product of matrices. We refer to features transforming according to (6) as $\rho_{\text{chan}}$-steerable, or features of type $\rho_{\text{chan}}$. This is a simpler form of the induced representations typically considered in steerable CNNs [14, 52], the simplification coming from the fact that we don’t enforce translation equivariance. Steerable ViT-features are illustrated in Figure 3.
@@
One can consider a patchified image as a steerable ViT feature in the following way. By patchification we mean the operation of reshaping an $M\times M\times C$ image first to $N^{2}$ patches of size $P\times P$, with $NP=M$ and then to an $N^{2}\times(P^{2}C)$ matrix. When we transform the original image by $\rho_{\text{image}}$, the patchified image is rotated by $\rho_{\text{token}}$ and $\rho_{\text{chan}}$ as in (6). Now $\rho_{\text{chan}}(g)$ is a permutation matrix that rotates or mirrors a patch and we will denote this particular $\rho_{\text{chan}}$ by $\rho_{\text{patch}}$.
We construct ViT versions that map images to steerable ViT features $\mathbf{x}\in\mathbb{R}^{L\times D}$ in the first layer (PatchEmbed) and then process steerable ViT features in subsequent layers. We choose $\rho_{\text{chan}}$ to be a multiple of $\rho_{\text{iso}}$, enabling efficient linear layers. Typically, we write $\rho_{\text{chan}}=\frac{D}{8}\rho_{\text{iso}}$ where $D$ is the embedding dimension of the ViT. For classification tasks, we map the steerable ViT features to $\mathrm{D}_{8}$ invariant features fed into a classification head. We denote ViTs that use $\rho_{\text{chan}}$-steerable features in all layers as $\mathrm{D}_{8}(\text{ViT})$. We will also consider networks that use $\rho_{\text{chan}}$-steerable features for the first layers and then map either to $D\rho_{\text{A1}}$-steerable features, these networks are denoted $\mathcal{I}_{8}(\text{ViT})$, or break equivariance, denoted $\mathcal{H}_{8}(\text{ViT})$. We refer to ViTs that fall into these three families broadly as octic ViTs. More details will be given in Section 4.2.

### v4

# Octic Vision Transformers: Quicker ViTs Through Equivariance

@@ Abstract
Why are state-of-the-art Vision Transformers (ViTs) not designed to exploit natural geometric symmetries such as 90-degree rotations and reflections? In this paper, we argue that there is no fundamental reason, and what has been missing is an efficient implementation. To this end, we introduce Octic Vision Transformers (octic ViTs) which rely on octic group equivariance to capture these symmetries. In contrast to prior equivariant models that increase computational cost, our octic linear layers achieve 5.33x reductions in FLOPs and up to 8x reductions in memory compared to ordinary linear layers. In full octic ViT blocks the computational reductions approach the reductions in the linear layers with increased embedding dimension. We study two new families of ViTs, built from octic blocks, that are either fully octic equivariant or break equivariance in the last part of the network. Training octic ViTs supervised (DeiT-III) and unsupervised (DINOv2) on ImageNet-1K, we find that they match baseline accuracy while at the same time providing substantial efficiency gains.
@@ 1 Introduction
In the pursuit of flexible yet scalable models, Vision Transformers (ViTs) (Dosovitskiy et al., 2021) have emerged as the dominant architecture in modern computer vision. Key to their success is the combination of visual tokens, constructed from image patches, with the powerful attention mechanism (Vaswani et al., 2017), resulting in a versatile and scalable architecture. This scalability is due in large part to weight-sharing between the tokens, which ensures permutation equivariance.
Equivariance provides a powerful inductive bias in neural networks by enforcing structured responses to transformations such as permutations, translations, rotations or reflections. Another major benefit of equivariance is the potential for reducing computational costs, for instance by token-wise weight-sharing as mentioned, or by parameterizing networks in the Fourier domain of a symmetry group. It is therefore interesting to investigate whether imbuing ViTs with further equivariance can yield improved compute efficiency. Equivariance to roto-reflections, formalized through the octic group $\mathrm{D}_{8}$, is a particularly attractive group, as many vision tasks exhibit such symmetries.
$\mathrm{D}_{8}$ equivariance was introduced to Convolutional Neural Networks (CNNs) by Cohen & Welling (2016), demonstrating improved parameter efficiency through weight sharing (Wood & Shawe-Taylor, 1996; Bekkers et al., 2018; Weiler & Cesa, 2019). Yet, despite its theoretical appeal, state-of-the-art vision models (Radford et al., 2021; Kirillov et al., 2023; Oquab et al., 2024; Wang et al., 2024a) do not incorporate group equivariance other than the permutation equivariance of transformer layers. We argue that this is not due to a lack of utility, but due to practical limitations: most existing implementations construct equivariant layers using computationally inefficient architectures, leading to increased FLOPs and runtime. Without efficient, hardware-compatible implementations, these methods have remained impractical for large-scale models. As a result, the potential of equivariant design remains largely unexplored in state-of-the-art systems, in particular in the context of ViTs. Recently, Bökman et al. (2025) showed that equipping ViTs with horizontal flip equivariance results in retained performance while saving FLOPs. However, their improvements are limited due to the small cardinality of their chosen group.
In this paper, we demonstrate that scaling equivariance to larger groups can be efficiently implemented – yielding faster, stronger, and more compact models, cf. Figure 1. Specifically, we introduce octic-equivariant layers for ViTs, leveraging the $\mathrm{D}_{8}$ symmetry group of $90^{\circ}$ rotations and reflections. Our approach integrates seamlessly into existing ViT architectures and leads to significant gains in throughput and memory efficiency, without sacrificing accuracy.
@@
    We introduce octic-equivariant layers for ViTs, described in Section˜3. In our implementation, octic-equivariant linear layers use $5.33$ times fewer FLOPs and $8$ times less memory per feature dimension than ordinary linear layers. Octic-equivariant ViT layers hence asymptotically have the same compute savings (Table˜1, Figure˜4(a)).
@@
    We propose two highly competitive families of ViTs, that are either fully rotation equivariant ($\mathcal{I}_{8}$) or break equivariance in late layers of the network ($\mathcal{H}_{8}$).
@@
    In Section˜4, we demonstrate empirically that our ViTs can be used in state-of-the-art ViT training recipes (DeiT III and DINOv2) without re-tuning hyperparameters. In particular, we achieve a $40\%$ FLOP saving with our $\mathcal{I}_{8}(\text{ViT-H})$ and $\mathcal{H}_{8}(\text{ViT-H})$ models while matching baseline performance (Figure˜1).
  4. (d)
    We study the effects of different methods of invariantization (testing six different methods in Appendix˜D), completely breaking equivariance (Section˜4.3) and varying the number of equivariant layers (Section˜4.3). These ablations will help guide future research on equivariant architectures at scale.
In addition to introducing new state-of-the-art ViTs, from a broader perspective our contributions give clear evidence that equivariance can matter at scale, which has been a subject of debate in recent literature (Abramson et al., 2024; Wang et al., 2024b; Brehmer et al., 2025; Bökman et al., 2025).
Figure 1: Computational savings. Using octic layers in ViTs significantly reduces the computational complexity without sacrificing accuracy on ImageNet-1K, for both supervised and self-supervised training. Detailed results can be found in Section 4.
@@ 2 Related Work
Vision Transformers.
The ViT was introduced by Dosovitskiy et al. (2021) and has subsequently achieved state-of-the-art results in many domains of computer vision (Carion et al., 2020; Radford et al., 2021; Kirillov et al., 2023; Edstedt et al., 2024; Wang et al., 2025). Significant efforts have been made to scale ViTs (Zhai et al., 2022; Dehghani et al., 2023), alongside strategies to do so efficiently (Alabdulmohsin et al., 2023). In this work, we propose incorporating octic layers in ViTs, which not only maintain efficiency at larger scales but become increasingly effective as model size grows, thus directly leveraging the scaling of ViTs. Hierarchical Transformers (Liu et al., 2021; Wu et al., 2021; Hassani et al., 2023) use translational symmetry, and SparseViT (Chen et al., 2023) extends this with sparse activations. In contrast, our work instead focuses on roto-reflections and builds exact equivariance into the architecture.
@@
The equivariance of CNNs to (cyclic) image translations can be extended to incorporate larger symmetry groups such as rotations and reflections, as shown by Cohen & Welling (2016); Dieleman et al. (2016) using Group Equivariant CNNs (G-CNNs). Cohen & Welling (2017); Weiler & Cesa (2019) generalized G-CNNs to steerable CNNs, where the features transform according to general group representations. Our octic ViTs can be seen as a more scalable ViT analogue of the octic steerable CNNs by Cohen & Welling (2017). There have also been prior efforts on attention- and Transformer-based equivariant architectures. For both point clouds (Fuchs et al., 2020; Hutchinson et al., 2021; Assaad et al., 2023; Liao & Smidt, 2023) and, more closely to ours, images (Romero et al., 2020; Xu et al., 2023; Rojas-Gomez et al., 2024; Kundu & Kondor, 2025). However, these prior works do not obtain computational benefits over non-equivariant transformers, in contrast to our ViTs. Our work is part of an ongoing research direction of studying and improving the scalability of equivariant networks (Bekkers et al., 2024; Brehmer et al., 2025; Bharadwaj et al., 2025; Vadgama et al., 2025; NVIDIA, ). Prior work in this direction mostly focused on point cloud data, with the notable exception of He et al. (2021) and Bökman et al. (2025) who considered images. He et al. (2021) increased the computational efficiency of G-CNNs but in contrast to our work did not achieve such benefits against standard networks. Directly inspiring our work, Bökman et al. (2025) demonstrated that incorporating horizontal mirroring equivariance into modern image classifiers increases compute efficiency while maintaining representational power. In contrast, we consider the larger octic group and thus achieve further savings in FLOPs. We conduct more extensive experimentation than Bökman et al. (2025) and address open questions in their work by studying the effect of breaking equivariance, invariantization and the number of equivariant blocks.
@@ 3 Method
In this section, we design octic-equivariant ViT layers. We begin with preliminaries for octic equivariance in Section 3.1, followed by the introduction of octic ViTs in Section 3.2, specifics of the Transformer layers in Section 3.3, and a detailed discussion of computational efficiency in Section 3.4. We summarize the most important notation in Section 3.1.1.
@@ 3.1 Preliminaries on Octic Equivariance
In this work we focus on the dihedral group with eight elements, $\mathrm{D}_{8}$ $\{e,r,r^{2},r^{3},s,sr,sr^{2},sr^{3}\}$, such that $r^{4}=s^{2}=e$ is the identity element and $r^{3}=srs$. [^1] The reader is cautioned that $\mathrm{D}_{8}$ is sometimes alternatively denoted $\mathrm{D}_{4}$. We think of $\mathrm{D}_{8}$ as acting on images by reflections $s$ and 90° rotations $r$. $\mathrm{D}_{8}$ is also called the octic group, and we opt for this shorter name throughout.
@@
[…]
We use the same notation as the original work on steerable CNNs (Cohen & Welling, 2017) for these irreps, but choose a different basis for $\rho_{E}$. It is known from elementary representation theory that any representation of $\mathrm{D}_{8}$ can be decomposed into irreps as
@@
so that $\rho_{\text{reg}}(g)$ can be written as a permutation matrix. Importantly, $\rho_{\text{reg}}$ commutes with pointwise activation functions such as GELU (Hendrycks & Gimpel, 2016).
@@
The regular representation $\rho_{\text{reg}}$ can be block-diagonalized to its isotypical decomposition $\rho_{\text{iso}}$ through equation 2 as $\rho_{\text{reg}}(g)=Q_{\text{reg}}\rho_{\text{iso}}(g)Q_{\text{reg}}^{-1}$ with
@@
It is a general fact for finite groups that the regular representation decomposes into all irreps with multiplicities equal to their dimensions. The change of basis $Q_{\text{reg}}$ (written out in full in Appendix˜A) is the inverse Fourier transform of $\mathrm{D}_{8}$, with $Q_{\text{reg}}^{-1}=Q_{\text{reg}}^{\mathsf{T}}$ being the Fourier transform.
[…]
Figure 2: $\mathrm{D}_{8}$ Linear layers. Implementing equivariant linear layers in the Fourier domain of $\mathrm{D}_{8}$ gives a major computational benefit. Left: A $C\times C$ weight matrix being multiplied by $L$ tokens of feature dimension $C$. Center: The block-diagonalization that happens when enforcing the layer to be $\mathrm{D}_{8}$-equivariant in the Fourier domain. More precisely, we enforce equivariance with respect to the representation $\frac{C}{8}\rho_{\text{iso}}$ that splits into irreps $\rho_{\text{A1}},\rho_{\text{A2}},\rho_{\text{B1}},\rho_{\text{B2}}$ and $\rho_{\text{E}}$ as detailed in Section 3. There is no mixing between different irreps and the weight sharing in the block-diagonal stems from the fact that $\rho_{\text{E}}$ is a two-dimensional irrep. Right: An efficient implementation of the original $C\times C$ by $C\times L$ matrix multiplication as four $\frac{C}{8}\times\frac{C}{8}$ by $\frac{C}{8}\times L$ and one $\frac{C}{4}\times\frac{C}{4}$ by $\frac{C}{4}\times 2L$ matrix multiplication. An equivariant linear layer of this type requires $16/3\approx 5.33$ times fewer FLOPs to compute and has 88 times fewer parameters than the ordinary linear layer shown to the left.
Equivariant networks use equivariant linear layers, which map between representations. A linear $G$-equivariant map or intertwiner, $W\in\mathbb{R}^{n\times n}$, between representations $\rho_{1},\rho_{2}$, commutes with their action, i.e., $\rho_{2}(g)W=W\rho_{1}(g)$. It follows from Schur’s lemma that $W=\lambda I$ for some scalar $\lambda$ if $\rho_{1},\rho_{2}$ are irreps, and that $\lambda=0$ if $\rho_{1},\rho_{2}$ are not isomorphic (Serre, 1977, Section I.2.2) [^2] We can apply Schur’s lemma for complex representations here since the irreps listed in Example 3.1 are irreducible over the complex numbers. We will however only use real-valued linear layers, i.e. $\lambda\in\mathbb{R}$.. As $\rho_{\text{iso}}$ is just a stack of irreducible representations, any intertwiner between such representations will be sparse, which is why they are less computationally expensive than ordinary linear layers. The naive computational complexity for a linear map $W:\mathbb{R}^{|\mathrm{D}_{8}|}\to\mathbb{R}^{|\mathrm{D}_{8}|}$ is $|\mathrm{D}_{8}|^{2}=64$ multiplications [^3] We ignore the additions for simplicity.. In contrast, intertwiners $\rho_{\text{iso}}\to\rho_{\text{iso}}$ require a total of $\sum_{i}^{k}m_{i}^{2}d_{i}=1+1+1+1+2^{2}\cdot 2=12$ multiplications for $\mathrm{D}_{8}$. From a signal processing perspective, this is analogous to the computational saving of convolution being point-wise multiplications in frequency space. We visualize the computational savings obtained by working in the Fourier domain of $\mathrm{D}_{8}$ in Figure 2.
@@
Square images can be considered as elements of $\mathbb{R}^{3\times M\times M}$ where $3$ is the number of color channels and $M$ is the image height/width in pixels. There is a natural $\mathrm{D}_{8}$-representation $\rho_{\text{image}}$ associated with square images, where $\rho_{\text{image}}(r)$ is the pixel permutation rotating the images anti-clockwise by 90° and $\rho_{\text{image}}(s)$ is the permutation reflecting the images left-to-right.
@@
In ViTs, features are elements of $\mathbb{R}^{C\times N\times N}$, which we will think of as $C\times N^{2}$ matrices. Here, $N$ is the number of image tokens along the height/width of the image, so $N=M/P$ where $P$ is the patch size (typically $P=14$ or $P=16$) and $C$ is the channel dimension. The simplest representation that we consider on features is the permutation representation $\rho_{\text{token}}$ that, analogously to $\rho_{\text{image}}$, permutes the tokens according to elements of $\mathrm{D}_{8}$.
@@
We can equip the channel dimension of ViT features with a group representation $\rho_{\text{chan}}$ to obtain “steerable” features. If $C$ is divisible by $8$, we can consider multiples of $\rho_{\text{reg}}$ or $\rho_{\text{iso}}$ as $\rho_{\text{chan}}$. The complete representation $\rho$ acting on the $C\times N^{2}$-matrix $\mathbf{x}$, is then permuting the tokens according to $\rho_{\text{token}}$ and modifying the channels according to $\rho_{\text{chan}}$. Concretely,
$\rho(g)\mathrm{Vec}(\mathbf{x})=\mathrm{Vec}\left(\rho_{\text{chan}}(g)\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}}\right)=\left(\rho_{\text{token}}(g)\otimes\rho_{\text{chan}}(g)\right)\mathrm{Vec}(\mathbf{x}),$ (6)
where $\mathrm{Vec}(\mathbf{x})$ is the column-wise vectorization of the matrix $\mathbf{x}$ and $\otimes$ is the tensor product of representations or (equivalently) the Kronecker product of matrices. We refer to features transforming according to equation 6 as $\rho_{\text{chan}}$-steerable, or features of type $\rho_{\text{chan}}$. This is a simpler form of the induced representations typically considered in steerable CNNs (Cohen & Welling, 2017; Weiler & Cesa, 2019), the simplification coming from the fact that we don’t enforce translation equivariance. Steerable ViT-features are illustrated in the Appendix, Figure 5.
@@
One can consider a patchified image as a steerable ViT feature in the following way. By patchification we mean the operation of reshaping a $3\times M\times M$ image first to $N^{2}$ patches of size $P\times P$, with $NP=M$ and then to a $3P^{2}\times N^{2}$ matrix. When we transform the original image by $\rho_{\text{image}}$, the patchified image is rotated by $\rho_{\text{token}}$ and $\rho_{\text{chan}}$ as in equation 6. Now $\rho_{\text{chan}}(g)$ is a permutation matrix that rotates or mirrors a patch and we will denote this particular $\rho_{\text{chan}}$ by $\rho_{\text{patch}}$.
3.1.1 Notation
For the convenience of the reader, we collect the most important notation in the paper in this section. We use the bold letter $\mathbf{x}$ for ViT features, which have shape $C\times L$, where $L$ is the number of tokens and $C$ the channel dimension. Typically, $L=N^{2}$, where $N$ is the image height/width in tokens, or $L=N^{2}+1$ with a class token. For an individual $C$-dimensional token we use the letter $x$ which is often acted on by $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$, in which case we can split $x$ into $C/8$-dimensional sub-tokens $x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E11}},x_{\text{E12}},x_{\text{E21}},x_{\text{E22}}$, where the first four transform according to the irreps $\rho_{\text{A1}},\rho_{\text{A2}},\rho_{\text{B1}},\rho_{\text{B2}}$ respectively while the $2\times\frac{C}{8}$-matrices $(x_{\text{E11}}\quad x_{\text{E12}})^{\mathsf{T}}$ and $(x_{\text{E21}}\quad x_{\text{E22}})^{\mathsf{T}}$ both transform according to $\rho_{\text{E}}$.
3.2 Octic ViTs
We construct ViT versions that map images to steerable ViT features $\mathbf{x}\in\mathbb{R}^{C\times L}$ in the first layer (PatchEmbed) and then process steerable ViT features in subsequent layers. We choose $\rho_{\text{chan}}$ to be a multiple of $\rho_{\text{iso}}$, enabling efficient linear layers. Typically, we write $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$ where $C$ is the embedding dimension of the ViT. For classification tasks, we map the steerable ViT features to $\mathrm{D}_{8}$ invariant features fed into a classification head. We denote ViTs that use $\rho_{\text{chan}}$-steerable features in all layers as $\mathrm{D}_{8}(\text{ViT})$. We will also consider networks that use $\rho_{\text{chan}}$-steerable features for the first layers and then map either to $D\rho_{\text{A1}}$-steerable features, these networks are denoted $\mathcal{I}_{8}(\text{ViT})$, or break equivariance, denoted $\mathcal{H}_{8}(\text{ViT})$. We refer to ViTs that fall into these three families broadly as octic ViTs. In Section 4.3, we study the effect of the number of octic layers.